OpenStudy (anonymous):

A Gaussian cube of side s is placed inside a uniformly charged insulating sphere of radius 2s and charge Q. What is the electric flux on the Gaussian cube? a. 3Q/(32pi Eo) b. 32Q/(3 pi Eo) c.) 3Q/(32 Eo) d.) 32Q/(3 Eo) e. Zero

OpenStudy (anonymous):

http://hyperphysics.phy-astr.gsu.edu/hbase/electric/gaulaw.html now u shud be able to find it

OpenStudy (anonymous):

The answer is a) 3Q/(32pi Eo) Solution: Since the sphere is insulating the charge is distributed uniformly throughout the volume. So the volume charge density= Q/volume of sphere = 3Q/4 pi (2s)^3 = 3Q/32 pi s^3. The flux through the cube = Charge enclosed by cube/Eo . Charge enclosed by cube = (volume charge density) * (volume of cube) = ( 3Q/32 pi s^3 ) * s^3 = 3Q/32 pi . => Flux = ( 3Q/32 pi )/Eo = 3Q/(32pi Eo)